OpenStudy (anonymous):

solve logbase5 (25x) its not 1/25 i tried it

OpenStudy (dcolley):

2

OpenStudy (anonymous):

Just remember this \[\log_a b^c = clog_a b\]

OpenStudy (anonymous):

if you have \[\log_5(25x)\] there is nothing to solve for. you would rewrite it as \[\log_5(25)+\log_5(x)=2+\log_5(x)\]

OpenStudy (anonymous):

unless it is \[\log_5(25^x)\] which is a different story

OpenStudy (anonymous):

@satellite73 ,Exactly, i just wanted to say that :D

OpenStudy (anonymous):

It was the 2+logbase5(x)

OpenStudy (anonymous):

can you help me do logbase3(9x)?

OpenStudy (anonymous):

exactly the same idea

OpenStudy (anonymous):

I see that you do logbase2 3+ logbase2 (x)= ?

OpenStudy (anonymous):

use \[\log(ab)=\log(a)+\log(b)\] to rewrite as \[\log_3(9)+\log_3(x)\]

OpenStudy (anonymous):

then note that \[\log_3(9)=2\] because \[3^2=9\]

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

so answer would be \[2+\log_3(x)\]

OpenStudy (anonymous):

ohhhhhhh ok thank you i was wondering where the 2 came from but now i see

OpenStudy (anonymous):

oh good. same place it came from in the first one right? \[\log_5(25)=2\] because \[5^2=25\]

OpenStudy (anonymous):

yes [=